OpenStudy (anonymous):

Factor: 2x2 + 7x + 3

OpenStudy (anonymous):

first you factorize so to factorize you multiply the first and the 3 number 2*3=6 make a list of factor 6 1*6 pick the factor that added make 7 2*3 so now you make a polynomial 2x^2 +6x +3x+ 3=0 now you make 2 grups of 2 factor so (2x^2 +6x) +(x+ 3) now you find the greatest common factor and divide by the parenthesis 2x(x+3)+1(x+3) the parenthesis have to be the same so now the factor outside the parenthesis are your other factors so (2x+1)(x+3) now you just set it equal to 0 so 2x+1=0 and solve 2x/2=-1/2 x=-1/2 x+3=3 x=-3

OpenStudy (anonymous):

(2x+1)(x+3)